Find the horizontal intercepts of the quadratic equation: 4x2 - 15x - 4 = 0.

Answer by Cromlix(4381)

4x^2 - 15x - 4 = 0
(4x + 1)(x - 4) = 0
4x + 1 = 0
x = -1/4
x - 4 = 0
x = 4
Horizontal intercepts are
(-1/4,0) and (4,0)
Hope this helps.
:-)